OpenStudy (anonymous):

use Lagrange multipliers to find the max and min of f(x,y) =x^2+y^2 subject to xy=1. Help please

OpenStudy (anonymous):

I got lamda = +-2. and stuck.

OpenStudy (amistre64):

what process got you to L=+-2 ?

OpenStudy (anonymous):

take del F = lamda *del g. it's <2x,2y> = lamda <y, x>

OpenStudy (anonymous):

so, 2x = lamda y --> x = lamda y/2 2y =lamda x ---> y = lamda x/2 time them xy =1 = lamda^2 /4 --> lamda = +-2

OpenStudy (amistre64):

Fx = 2x , LGx = Ly x = Ly/2 Fy = 2y , LGy = Lx y = Lx/2 and we know that xy = 1 L^2 xy/4 = 1 L^2 xy = 4 since xy=1, L = +- 2 does seem appropriate :) so far

OpenStudy (anonymous):

and then? I stuck.

OpenStudy (amistre64):

x = y; y = x or x = -y; y = -x would be our possible outcomes for L = +- 2 if x=y then we can say x^2 = 1 if y=-x , does -x^2 = 1 exist?

OpenStudy (anonymous):

how x =-y? I got x =y but x cannot be -y

OpenStudy (amistre64):